OpenStudy (mrhoola):

A 10 gram bullet hits and becomes embedded in a 2 kg block of wood which is initially at rest. What was the bullet's speed before the collision if the combination has a speed of 3 m/s after the collision? Linear momentum : my question is - Is this an isolated system or non-isolated? Please tell me which one .

OpenStudy (p0sitr0n):

need to do it with forces or energy?

OpenStudy (mrhoola):

Energy

OpenStudy (3psilon):

So use conservation of linear momentum p = mv so \[m_{1}v + m_{2}v = (m_{1}+m_{2})V_{f}\] Now plug in .The wood is at rest so v wood = 0 \[m_{1}v + m_{2}(0) = (m_{1}+m_{2})3\] 10(v) = 12(3) 36/10 = v initial of bullet

OpenStudy (mrhoola):

Thank you !